# FLawN-T5: An Empirical Examination of Effective Instruction-Tuning Data Mixtures for Legal Reasoning

### v1

@@ B.2 Evaluation
B.2.1 LegalBench Dataset Held Out
If the source dataset of the LegalBench task occurs in LawInstruct, we remove it from the evaluation. Below, we list which tasks are overlapping. Overall 100 tasks are held out, so 61 tasks are remaining for LegalBench evaluation.
  • ContractNLI
    • LawInstruct: ContractNLI-contract_nli
    • LegalBench: contract_nli_*
  • CUAD
    • LawInstruct: NaturalInstructionsLegal-cuad_answer_generation, NaturalInstructionsLegal-cuad_question_generation
    • LegalBench: cuad_*
  • GLOBALCIT Citizenship Law Dataset
    • LawInstruct: InternationalCitizenshipLawQuestions-international_citizenship_law_questions_mode_acq, InternationalCitizenshipLawQuestions-international_citizenship_law_questions_mode_loss
    • LegalBench: international_citizenship_questions
  • MAUD
    • LawInstruct: MAUD-answer, MAUD-category, MAUD-question, MAUD-text_type
    • LegalBench: maud_*
  • OPP-115 (Online Privacy Policies, set of 115) Corpus
    • LawInstruct: NaturalInstructionsLegal-online_privacy_policy_text_information_type_generation, NaturalInstructionsLegal-online_privacy_policy_text_purpose_answer_generation
    • LegalBench: opp_115_*
  • Overruling
    • LawInstruct: NaturalInstructionsLegal-overruling_legal_classification
    • LegalBench: overruling
  • PrivacyQA
    • LawInstruct: PrivacyQA-privacy_qa
    • LegalBench: privacy_policy_qa
@@ B.2.1 LegalBench Dataset Held Out
  • StAtutory Reasoning Assessment (SARA)
    • LawInstruct: Sara-sara_entailment, Sara-sara_tax_liability, SaraProlog-sara_prolog_facts, SaraProlog-sara_prolog_statute
    • LegalBench: sara_* (built off of SARA v2)
  • Unfair Terms of Service
    • LawInstruct: This dataset is duplicated between LexGLUE-unfair_tos, LEXTREME-online_terms_of_service_clause_topics (multilingual version), LEXTREME-online_terms_of_service_unfairness_levels (multilingual version)
    • LegalBench: unfair_tos
B.2.2 LegalBench Task Held Out
We additionally catalog instructions which train the LLM for a task captured in LegalBench. It is not necessary that the instruction-response pair in LawInstruct contain data from LegalBench, just that they are about similar legal tasks (e.g., classifying choice-of-forum provisions). Below, we list which tasks are overlapping. Overall 64 tasks are held out, so 97 tasks are remaining for LegalBench evaluation.
  • Rhetorical Role Labeling (Evaluates LLM’s ability to annotate the function of a piece of legal text.)
    • LawInstruct: bva_decisions_label, indian_text_segmentation, german_argument_mining
    • LegalBench: function_of_decision_section, oral_argument_question_purpose
  • Civil Procedure Questions (Evaluates LLM’s application of basic laws of US jurisdiction.)
    • LawInstruct: civipro_questions_generate_*
    • LegalBench: diversity_*, personal_jurisdiction
  • Legal Entailment (Evaluates LLM’s understanding of logical relationships in legal texts.)
    • LawInstruct: coliee_task3_passage_entailment, contract_nli, lawng_nli_entailment
    • LegalBench: contract_nli_*
  • Contractual Clause Classification (Evaluates LLM’s ability to classify contractual clauses.)
    • LawInstruct: unfair_tos, german_rental_agreements
    • LegalBench: cuad_*, jcrew_blocker, unfair_tos, contract_qa
Appendix C Additional Ablations
C.1 Crosslingual Transfer from Multilingual Data
@@
Is there crosslingual transfer from multilingual data? $\Rightarrow$ On the English LegalBench, we do not see any crosslingual transfer. In Figure 13, we compare the performance of the complete multilingual instruction dataset and the English subset across two differently licensed datasets (research and commercial licenses). We see no statistically significant difference between the multilingual training and the English training. We also see no difference between the differently licensed datasets. This means that just training on the commercial subset is enough. We show detailed results on individual LegalBench categories in Appendix D Table 7. Per default we use the English dataset in all following experiments unless specified otherwise.
Appendix D Detailed Results
@@
Table 10: Flan-T5 Small models with different domain adaptation strategies (amount of IFT data during continued pretraining).
@@
Table 11: Flan-T5 Base models with different domain adaptation strategies (amount of IFT data during continued pretraining).
@@
Table 12: Flan-T5 XL models with different domain adaptation strategies (amount of IFT data during continued pretraining).
@@
Table 13: Flan-T5 XXL models with different domain adaptation strategies (amount of IFT data during continued pretraining).

### v2

@@ C.2 Evaluation
C.2.1 LegalBench Dataset Held Out
If the source dataset of the LegalBench task occurs in LawInstruct, we remove it from the evaluation. Below, we list which tasks are overlapping. Overall 100 tasks are held out (see Table 3 for the complete list), so 61 tasks are remaining for LegalBench evaluation.
Table 3: LegalBench Dataset Held Out
Dataset LawInstruct LegalBench
ContractNLI ContractNLI-contract_nli contract_nli_*
CUAD NaturalInstructionsLegal-cuad_answer_generation, NaturalInstructionsLegal-cuad_question_generation cuad_*
GLOBALCIT Citizenship Law Dataset InternationalCitizenshipLawQuestions-international_citizenship_law_questions_mode_acq, InternationalCitizenshipLawQuestions-international_citizenship_law_questions_mode_loss international_citizenship_questions
MAUD MAUD-answer, MAUD-category, MAUD-question, MAUD-text_type maud_*
OPP-115 (Online Privacy Policies, set of 115) Corpus NaturalInstructionsLegal-online_privacy_policy_text_information_type_generation, NaturalInstructionsLegal-online_privacy_policy_text_purpose_answer_generation opp_115_*
Overruling NaturalInstructionsLegal-overruling_legal_classification overruling
PrivacyQA PrivacyQA-privacy_qa privacy_policy_qa
@@ C.2.1 LegalBench Dataset Held Out
StAtutory Reasoning Assessment (SARA) Sara-sara_entailment, Sara-sara_tax_liability, SaraProlog-sara_prolog_facts, SaraProlog-sara_prolog_statute sara_* (built off of SARA v2)
Unfair Terms of Service LexGLUE-unfair_tos, LEXTREME-online_terms_of_service_clause_topics (multilingual version), LEXTREME-online_terms_of_service_unfairness_levels (multilingual version) unfair_tos
C.2.2 LegalBench Task Held Out
We additionally catalog instructions which train the LLM for a task captured in LegalBench. It is not necessary that the instruction-response pair in LawInstruct contain data from LegalBench, just that they are about similar legal tasks (e.g., classifying choice-of-forum provisions). In Table 4, we list which tasks are overlapping. Overall 64 tasks are held out, so 97 tasks are remaining for LegalBench evaluation.
Table 4: LegalBench Task Held Out
Task LawInstruct LegalBench
Rhetorical Role Labeling bva_decisions_label, indian_text_segmentation, german_argument_mining function_of_decision_section, oral_argument_question_purpose
Civil Procedure Questions civipro_questions_generate_* diversity_*, personal_jurisdiction
Legal Entailment coliee_task3_passage_entailment, contract_nli, lawng_nli_entailment contract_nli_*
Contractual Clause Classification unfair_tos, german_rental_agreements cuad_*, jcrew_blocker, unfair_tos, contract_qa
Appendix D Additional Ablations
D.1 Sampling Style
Should we sample each dataset equally or rather by the number of examples? $\Rightarrow$ Sampling by the number of examples generally leads to better performance. In Figure 12, we compare the performance of two sampling styles (equal sampling of each dataset and sampling by the number of examples) across both the research and commercial licensed dataset (detailed results in Appendix E Table 8). For the XL and XXL sizes, sampling by the number of examples is better than equal weight for datasets for both the research and commercial datasets, although not always statistically significant (XL res $p=0.049$, XL comm $p=0.052$, XXL res $p<0.001$, XXL comm $p=0.31$). For the Small size, sampling by the number of examples is better for the research dataset ($p<0.001$) but not for the commercial dataset ($p=0.099$), while there is no difference for the Base size. By default, we sample by the number of examples in all following experiments unless specified otherwise.
D.2 License of Instruction Tuning Datasets
Do we need data licensed non-commercially for good performance? $\Rightarrow$ The commercially licensed data seems to be enough for the larger models. In Figure 12, we compare the performance of two differently licensed datasets (research and commercial licenses) across both sampling each dataset equally and by the number of examples (detailed results in Appendix E Table 8). There are fewer datasets available with more permissive licenses allowing for commercial use than for research use (see Table 2 for details on licenses). Except for Small size ($p<0.001$), using more diverse data available only for research shows no significant benefit. By default, we use the commercially licensed dataset in all subsequent experiments unless specified.
D.3 Crosslingual Transfer from Multilingual Data
@@
Is there crosslingual transfer from multilingual data? $\Rightarrow$ On the English LegalBench, we do not see any crosslingual transfer. In Figure 13, we compare the performance of the complete multilingual instruction dataset and the English subset across two differently licensed datasets (research and commercial licenses). We see no statistically significant difference between the multilingual training and the English training. We also see no difference between the differently licensed datasets. This means that just training on the commercial subset is enough. We show detailed results on individual LegalBench categories in Appendix E Table 9. Per default we use the English dataset in all following experiments unless specified otherwise.
Appendix E Detailed Results
@@
Table 12: Flan-T5 Small models with different domain adaptation strategies (amount of IFT data during continued pretraining). 1-IFT-to-X-PRE means that for every X pretraining examples we mix in one instruction example. ONLY-PRE means we did not mix in any instruction examples.
@@
Table 13: Flan-T5 Base models with different domain adaptation strategies (amount of IFT data during continued pretraining). 1-IFT-to-X-PRE means that for every X pretraining examples we mix in one instruction example. ONLY-PRE means we did not mix in any instruction examples.
@@
Table 14: Flan-T5 XL models with different domain adaptation strategies (amount of IFT data during continued pretraining). 1-IFT-to-X-PRE means that for every X pretraining examples we mix in one instruction example. ONLY-PRE means we did not mix in any instruction examples.
@@
Table 15: Flan-T5 XXL models with different domain adaptation strategies (amount of IFT data during continued pretraining). 1-IFT-to-X-PRE means that for every X pretraining examples we mix in one instruction example. ONLY-PRE means we did not mix in any instruction examples.